Michael is one-fourth as old as his father. Next year, their ages will total 42 years. How old is each now?
Answer by macston(5194)

M=Michael's age; F=father's age=4M; F+M=40 (42-2=Sum of ages this year=40)
Substitute F=4M in F+M=40
4M+M=40
5M=40
M=8 ANSWER Michael is 8 years old
F=4M=4(8)=32 ANSWER Father is 32 years old
CHECK
Next year sum of ages=42
(F+1)+(M+1)=42 years
(32+1yr)+(8+1yr)=42 yrs
33yrs+9yrs=42 yrs
42 yrs=42 yrs